the average human being contains 60 trillion living cells. if there are currently 7.217 million people on earth. how many moles of living human cells are currently on our planet? if 150billion humans have inhabited the planet in all of history, how many moles of human cells have existed so far on our planet?
1
Expert's answer
2020-07-27T05:32:50-0400

As one human contains 60 trillion cells and there are 7,217 million people, the total number of living cells equals:

Ncells = 60 × 1012 × 7.217 × 109 = 4.33 × 1023

The number of moles can be calculated by the following formula:

n = Ncells / NA

where n - number of moles, Ncells - total number of cells, NA - Avogadro's number (6.022 × 1023).

ncells = 4.33 × 1023 / 6.022 × 1023 = 0.72 mol.

In the case of 150 billion humans:

Ncells = 60 × 1012 × 150 × 109 = 9 × 1024

ncells = 9 × 1024 / 6.022 × 1023 = 15 mol


Answer: 0.72 mol and 15 mol.